If a freight carrier charges 55 cents for a package up to one ounce and 39 cents for each additional ounce or fraction of an ounce. Find the cost of shipping a package that weighs 168 grams.
A) $2.50 B) $3.69 C) $2.89 D) $2.73

Answer by jim_thompson5910(35256)

Let's use the idea that 1 ounce is roughly 28.3495231 grams to determine how many ounces are in 168 grams.




















-------------------------------------------------------------------------------


Since 168 grams is roughly equal to 5.92602528 ounces, this means that we


Step 1) Get charged 55 cents for the first ounce


Step 2) Get charged 39 cents, per ounce, for the additional 5 ounces (round 5.92602528 to 6, then note how 6-1 = 5).


So following those steps above we have 55+5*39 = 250 as the total cost in cents


Divide that by 100 to convert to dollars: 250/100 = 2.50 so that points to the answer being choice A